OpenStudy (twizttiez):

Find the constant of variation for the relationship shown in the following table: x 1 2 3 4 y 4 8 12 16 A. 1 B. 2 C. 3 D. 4

OpenStudy (twizttiez):

I think it's D @Data_LG2

OpenStudy (twizttiez):

Sorry that I tagged you i just need some help

OpenStudy (anonymous):

:) can you tell me why you choose D? no worries, i'm willing to help as long as I know the topic :D

OpenStudy (twizttiez):

D.4 goes in a pattern and it's the only one that looks to make sense is why i chose D lol

OpenStudy (twizttiez):

Now whether or not that made sense idk lol

OpenStudy (anonymous):

yes that's right! constant variation is the ratio between two variables in a direct variation or the product of two variables in an inverse variation. ~google

OpenStudy (twizttiez):

YAY!! can you go and see if i was right on another ? idk if i was or not

OpenStudy (anonymous):

sure, post it here wait, is this a quiz or some kind of test? (just want to make sure)

OpenStudy (twizttiez):

No just a jumble of questions

OpenStudy (anonymous):

alright (:

OpenStudy (twizttiez):

It's in my book i just need to know if i got it down ya know?

OpenStudy (anonymous):

kk ^_^

OpenStudy (twizttiez):

Which equations show a direct variation between x and y when solved for y? Choose all answers that are correct. A. -10+y=3x-10 B. y=-1/x C. y+9=9x D. y=x-1 I think my answers are A,C and D

OpenStudy (twizttiez):

@Data_LG2